5x + 3y= 900 and 3x + 5y=860. Find the value of x and y
Answer by Alan3354(69443)

5x + 3y= 900
3x + 5y= 860
---------
25x+ 15y = 4500
9x + 15y = 2580
--------------------- Subtract